two numbershave awhat is the smaller of the two numbers sum of 91 their difference is 3
what is the smaller of the two numbers

Answer by oberobic(2304)

Assuming...
Sum of two numbers is 91.
We can call the two numbers x & y.
x + y = 91
x = 91-y
.
The difference of the two numbers is 3, which we can assume to be:
x - y = 3
x = y+3
.
x + y = 91
.
Substitute for x.
y+3 +y = 91
Collect
2y +3 = 91
Subtract 3 from both sides
2y = 88
Divide both sides by 2
y = 44
.
Since x = y+3,
x = 47
.
Answer:
The two numbers are 44 and 47.
.
Done.